# Estimating accretion times of halo substructures in the Milky Way

### v1

@@ 4.1 Estimating the time of accretion
gomez10Identificationa and gomez10Identification proposed a method for estimating accretion times from orbital frequencies and validated it using simulated data. Although they were unable to apply the method to real observations available then due to its stringent precision requirements, the high-precision data from Gaia DR3 now provide the necessary foundation to do so. We adjust some details of the method and describe the full procedure below.
@@
In step 1, we compute the bi-dimensional histogram in the $\Omega_{r}-\Omega_{\phi}$ plane individually for each substructure, with a bin width of $\Delta$ in both dimensions. The bin width $\Delta$ may differ among substructures and will be described later. For both dimensions, the number of bins $N$ is required to be equal, odd, and $\geqslant 201$.
@@
$P(k_{r})=\begin{cases}\frac{1}{N^{2}}\sum_{k_{\phi}=-k_{\mathrm{slit}}}^{k_{\mathrm{slit}}}|H(0,k_{\phi})|^{2}&,k_{r}=0,\\ \frac{1}{N^{2}}\sum_{k_{\phi}=-k_{\mathrm{slit}}}^{k_{\mathrm{slit}}}\left(|H(k_{r},k_{\phi})|^{2}+|H(-k_{r},k_{\phi})|^{2}\right)&,k_{r}>0,\end{cases}$
@@
4.2 Significances and uncertainties
To estimate the significances and uncertainties of the accretion time results, the above method is applied to the 1,000 realizations of both the substructures and the random samples. For the substructures, we only consider member stars with small orbital frequency uncertainties, with thresholds that vary among substructures and are specified below.
@@ 4.2 Significances and uncertainties
Figure 5: Median 1-D power spectra for substructure GL-1 (black solid lines) and random samples (blue solid lines). The $1\sigma$ and $3\sigma$ intervals of the random samples are shown as dark blue and light blue shaded regions, respectively. The $x$-axes represent $2\pi$ times the wavenumbers in the $r$ (top panels) and $\phi$ (bottom panels) directions. The estimated accretion times are marked by black vertical dashed lines, with significances given in the legend. Left and right panels show the results using the eilers19Circular and bovy15Galpy potentials, respectively. Subtitles indicate the substructure name, number of stars, bin width $\Delta$, and applied uncertainty selection.
Using the substructure GL-1 as an example, we adopt a bin width $\Delta$ of 0.3 and select member stars with $\sigma_{\Omega_{r}}<0.2$ Gyr-1 and $\sigma_{\Omega_{\phi}}<0.2$ Gyr-1 to estimate the accretion time. A total of 1,000 sets of 1-D power spectra $P(k_{r})$ and $P(k_{\phi})$ are computed, but the corresponding accretion times are not calculated at this stage. The black solid lines on the left panels of Figure 5 show the medians of the power spectra as functions of $2\pi$ times the wavenumber. With this choice of the $x$-axis, the accretion time can be obtained directly. The black vertical dashed lines mark the locations of the maximum peaks, corresponding to accretion times of 5.6 Gyr and 5.8 Gyr for the $r$ and $\phi$ directions, respectively.
Applying the above selection on orbital frequency uncertainties results in a sample of 111 stars. For comparison, we randomly draw an equal number of stars from the halo sample 1,000 times. We then compute the 1-D power spectra of these samples, and the medians of the spectra are indicated by the blue lines in the left panels of Figure 5. The dark blue and light blue shaded regions represent the $1\sigma$ and $3\sigma$ intervals, respectively. It can be seen that, at larger wavenumbers, the power spectra are nearly constant, corresponding to contributions from random noise.
@@
Figure 6: Histograms of the $r$ (left) and $\phi$ (right) direction power spectra in the high wavenumber region. The blue curves represent the fitted models of the two-parameter gamma distribution. The black vertical lines mark the maximum peak of the median power spectra of substructure GL-1.
@@
The fitting results are shown in Figure 6. The significance of the highest peak in the median power spectrum of the substructure is estimated using the cumulative distribution function, and the peak is indicated by the vertical black line in the figure. In the case of GL-1, the accretion times in the $r$ and $\phi$ directions have significances of $4.4\sigma$ and $4.5\sigma$, respectively.
@@
Uncertainties in the accretion time are estimated from the 1,000 sets of 1-D power spectra associated with the realizations. Due to the influence of errors, the highest peaks of the individual 1-D power spectra may not coincide with that of the median power spectrum. Therefore, we first estimate the range of accretion times using the median power spectrum, and then identify the accretion time of each individual power spectrum within this range.
@@
To obtain this range, the median power spectrum of GL-1 is smoothed using gaussian_filter1d of SciPy (virtanen20SciPy). We adopt a full width at half maximum of 0.2 Gyr, corresponding to $0.2/2\pi\approx 0.03$ Gyr in wavenumber space. The upper and lower limits of the range are defined as the nearest troughs on either side of the highest peak, if the median of the power spectra of the troughs is lower than that of the random samples. If no trough on the left side of the peak satisfies the condition, the minimum on that side is taken as a substitute. In the analogous case on the right side, the upper bound is taken to be infinite.
@@
For each individual power spectrum, the accretion time is determined within this range, with the procedure applied separately for the $r$ and $\phi$ directions. The accretion time of GL-1 is estimated to be $5.55\pm 0.08$ Gyr from the $r$ direction and $5.84\pm 0.20$ Gyr from the $\phi$ direction. The resulting combined accretion time is $5.59^{+0.07}_{-0.08}$ Gyr.
@@
The bin width $\Delta$ and the uncertainty threshold are determined by a grid search to maximize the significance. Their values are taken from $e^{-2}\approx 0.14$ to $e\approx 2.72$, with steps of 0.1 in log space. We only consider results with more than 20 stars after the uncertainty selection.
4.3 Effect of different potential
@@
Figure 7: Circular velocity curves $v_{c}$ of the eilers19Circular and bovy15Galpy potentials as a function of Galactocentric radius $R$.
@@
The estimated accretion times vary with the choice of the Milky Way potential model. For comparison, we adopt the potential of bovy15Galpy. The circular velocity curves of this potential and the eilers19Circular potential used above are shown in Figure 7. As shown in this figure, the B15 potential has a smaller mass than the E19. At the solar position, the circular velocity is 219 km s-1 for the B15 potential and 229 km s-1 for the E19 potential.
@@
The bin width $\Delta$ and the uncertainty threshold are adjusted to maximize the significance. The right panels of Figure 5 show the 1-D power spectra of GL-1 computed using the B15 potential. It can be seen that the significances are noticeably reduced, especially for the $r$ direction. The estimated accretion times also differ, but the results from the two potentials are not entirely inconsistent. In the 1-D power spectra based on the B15 potential, a peak can still be identified around 5.7 Gyr, corresponding to the result from the E19 potential, although it is no longer the most prominent one.
@@
4.4 Accretion times of other substructures
@@
Figure 9: The differences in orbital frequencies between the eilers19Circular and the bovy15Galpy potentials ($\Omega_{\mathrm{Bovy}}-\Omega_{\mathrm{Eilers}}$) for member stars of different substructures versus orbital pericenters under the eilers19Circular potential. Different substructures are indicated by different colors. Left and right panels show the frequency differences in the $r$ and $\phi$ directions, respectively.
For GL-4, the accretion time inferred from the $\phi$ direction analysis under the B15 potential is $4.0\pm 0.2$ Gyr, with a significance of $3.3\sigma$. For the E19 potential, the result is similar, with an accretion time of $3.9_{-0.26}^{+0.36}$ Gyr and a lower significance of $1.7\sigma$. The higher significance of the result under the B15 potential may be due to the pericenter of GL-4 being closer to the Milky Way center. Figure 9 shows the distribution of the differences in orbital frequencies between the two potentials, $\Delta_{\Omega}=\Omega_{\mathrm{Bovy}}-\Omega_{\mathrm{Eilers}}$, versus the orbital pericenters in the E19 potential. Member stars of different substructures are shown in different colors. A negative $\Delta\Omega$ indicates that the orbital frequency is smaller in the B15 potential.
Furthermore, this difference can also provide guidance for improving the potential. eilers19Circular adopted the same baryonic components as pouliasis17Milky and fitted the dark matter halo using the circular velocity curve within $5\leqslant R\leqslant 25$ kpc. The region $R<5$ kpc lies outside the fitted domain, and therefore the E19 potential is unreliable there. Since the result based on the B15 potential shows higher significance, the E19 potential may overestimate the mass within $R\sim 5$ kpc.
@@
A similar situation occurs for GR-1, as shown in Figure 9, the pericenters of its member stars are closer to the Milky Way center compared to those of GL-4, which is also the region where the E19 and B15 potentials exhibit the largest discrepancy. Its 1-D power spectra are shown in Figure 10. For the $r$ direction, the results from the two potentials are $2.2\pm 0.1$ and $2.3\pm 0.1$ Gyr, which are fairly close. For the E19 potential, the significance is $3.0\sigma$, whereas for the B15 potential, it is higher, at $3.7\sigma$.
For the E19 potential, the power spectrum in the $\phi$ direction shows a trough near 2.2 Gyr, even lying outside the $3\sigma$ interval. The inconsistency of the E19 potential results between the two directions suggests a limitation of the potential. Under the B15 potential, a small peak appears in the $\phi$ direction power spectrum near 2.3 Gyr, which hints that the E19 potential may overestimate the rotation curve for $R<5$ kpc. This is consistent with the inference based on the GL-4 results.
The pericenters of the member stars of GR-1 are very close to the Milky Way center, where the potential is dominated by the bulge. The comparison between the E19 and B15 potential results may suggest a less massive bulge, i.e., with a mass smaller than the $1.1\times 10^{10}$ $M_{\odot}$ of the bulge in the E19 potential. For the B15 potential with a bulge mass of $0.5\times 10^{10}$ $M_{\odot}$, the peak around 2.3 Gyr in the $\phi$ direction is also very low, indicating that the bulge is likewise questionable.
@@
For the substructure GL-1, the accretion time results based on the orbital frequencies in the $r$ and $\phi$ directions show good consistency, yielding $5.55\pm 0.08$ Gyr and $5.84\pm 0.20$ Gyr, respectively. The statistical significance is also high, at $4.4\sigma$ and $4.5\sigma$, respectively. The agreement between the two directions and their high significance indicate the robustness of the results. The accretion time for GL-1, obtained by combining the $r$ and $\phi$ direction estimates, is $5.59^{+0.07}_{-0.08}$ Gyr. When the B15 potential is used instead of the E19 potential, which is based on the more reliable Gaia data, the results change and the significances decrease substantially, indicating that the results are sensitive to the choice of potential.
For GL-4 the inferred accretion time is $4.0\pm 0.2$ Gyr with a significance of $3.3\sigma$, while for GR-1 it is $2.3\pm 0.1$ Gyr with a significance of $3.7\sigma$. They exhibit high significance only in a single direction, and only under the B15 potential. This may be because their member stars have pericenters closer to the Milky Way center, where the E19 potential is extrapolated beyond its fitted range and thus overestimates the mass at $R<5$ kpc. For the substructure GR-1, which lies closer to the Milky Way center, the comparison of the $r$ and $\phi$ direction estimates under the E19 potential suggests that the bulge mass may be smaller than the E19 value of $1.1\times 10^{10}$ $M_{\odot}$. For further refining accretion times for these and other phase-space substructures to obtain the accretion history of the Milky Way, more precise astrometric data provided by, e.g., Gaia DR4, China Space Station Survey Telescope and Roman Space Telescope, are needed.
This work is supported by the National Key Research and Development Program of China No. 2024YFA1611902, National Natural Science Foundation of China (NSFC) No. 12588202, CAS Project for Young Scientists in Basic Research grant No. YSBR-062, the Strategic Priority Research Program of Chinese Academy of Sciences grant No. XDB1160102, and the science research grants from the China Manned Space Project with NO. CMS-CSST-2025-A11. M.C. acknowledges support in part from JSPS KAKENHI (No. JP24K00669 and 25H00394). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.

### v2

@@ IV.1 Estimating the time of accretion
Although they were unable to apply the method to real observations available then due to its stringent precision requirements, the high-precision data from Gaia DR3 now provide the necessary foundation to do so. We adjust some details of the method and describe the full procedure below.
@@
In step 1, we compute the bi-dimensional histogram in the $\Omega_{r}-\Omega_{\phi}$ plane individually for each substructure, with a bin width of $\Delta$ in both dimensions. The bin width $\Delta$ may differ among substructures and will be described later. For both dimensions, the number of bins $N$ is required to be equal, odd, and $\geqslant 201$. To reduce sensitivity to the choice of bin edge positions, we compute histograms for each dimension with the bin edges shifted by 0, 0.2, 0.4, 0.6, and 0.8 times the bin width $\Delta$, and constructed all combinations of these shifts across the two dimensions.
@@
$P(k_{r})=\begin{cases}\dfrac{1}{N^{2}}\displaystyle\sum_{k_{\phi}=-k_{\mathrm{slit}}}^{k_{\mathrm{slit}}}|H(0,k_{\phi})|^{2},\\ \hfill\mathrm{if}\;k_{r}=0,\\[6.00006pt] \dfrac{1}{N^{2}}\displaystyle\sum_{k_{\phi}=-k_{\mathrm{slit}}}^{k_{\mathrm{slit}}}\left(|H(k_{r},k_{\phi})|^{2}+|H(-k_{r},k_{\phi})|^{2}\right),\\ \hfill\mathrm{if}\;k_{r}>0,\end{cases}$ (5)
@@
IV.2 Significances and uncertainties
To estimate the significances and uncertainties of the accretion time results, the above method is applied to the 1,000 realizations of both the substructures and the random samples. For the substructures, we only consider member stars whose orbital frequencies have small uncertainties, with thresholds varying among substructures as specified below.
@@ IV.2 Significances and uncertainties
Figure 5: Median 1-D power spectra for substructure GL-1 (black solid curves) and random samples (blue dashed curves). The $1\sigma$ and $3\sigma$ intervals of the random samples are shown as dark blue and light blue shaded regions, respectively. The $x$-axes represent $2\pi$ times the wavenumbers in the $r$ (top panels) and $\phi$ (bottom panels) directions. The estimated accretion times are marked by black vertical dashed lines, with significances given in the legend. Left and right panels show the results using the Eilers et al. (2019) and Bovy (2015) potentials, respectively. Subtitles indicate the substructure name, number of stars, bin width $\Delta$, and applied uncertainty selection.
Using the substructure GL-1 as an example, we adopt a bin width $\Delta$ of 0.41 and select member stars with $\sigma_{\Omega_{r}}<0.15$ Gyr-1 and $\sigma_{\Omega_{\phi}}<0.15$ Gyr-1 to estimate the accretion time. The choice of the bin edge shifts will be discussed later. A total of 1,000 sets of 1-D power spectra $P(k_{r})$ and $P(k_{\phi})$ are computed, but the corresponding accretion times are not calculated at this stage. The black solid curves on the left panels of Figure 5 show the medians of the power spectra as functions of $2\pi$ times the wavenumber. With this choice of the $x$-axis, the accretion time can be obtained directly. The black vertical dashed lines mark the locations of the maximum peaks, corresponding to accretion times of 5.6 Gyr and 5.7 Gyr for the $r$ and $\phi$ directions, respectively.
Applying the above selection on orbital frequency uncertainties results in a sample of 89 stars. For comparison, we randomly draw an equal number of stars from the halo sample 1,000 times. We then compute the 1-D power spectra of these samples, and the medians of the spectra are indicated by the blue dashed curves in the left panels of Figure 5. The dark blue and light blue shaded regions represent the $1\sigma$ and $3\sigma$ intervals, respectively. It can be seen that, at larger wavenumbers, the power spectra are nearly constant, corresponding to contributions from random noise.
@@
Figure 6: Histograms of the $r$ (left) and $\phi$ (right) direction power spectra in the high wavenumber region. The dashed lines represent the fitted models of the two-parameter gamma distribution. The vertical solid lines mark the maximum peak of the median power spectra of substructure GL-1.
@@
The fitting results are shown in Figure 6. The significance of the highest peak in the median power spectrum of the substructure is estimated using the cumulative distribution function, and the peak is indicated by the black vertical line in the figure. In the case of GL-1, the accretion times in the $r$ and $\phi$ directions have significances of $4.3\sigma$ and $3.9\sigma$, respectively.
@@
Uncertainties in the accretion time are estimated from the 1,000 sets of 1-D power spectra associated with the realizations. Due to errors, the peak in each realization’s 1-D power spectrum that near the highest peak of the median power spectrum is not always the highest peak. Therefore, we adopt a prior that constrains each realization’s accretion time estimate to lie around the value derived from the median power spectrum. This prior is implemented as a uniform distribution over a finite interval.
@@
To obtain this interval, the median power spectrum of GL-1 is smoothed using gaussian_filter1d of SciPy (Virtanen et al., 2020). We adopt a full width at half maximum of 0.2 Gyr, corresponding to $0.2/2\pi\approx 0.03$ Gyr in wavenumber space. The upper and lower limits of the interval are defined as the nearest troughs on either side of the highest peak, if the median of the power spectra of the troughs is lower than that of the random samples. If no trough on the left side of the peak satisfies the condition, the minimum on that side is taken as a substitute. In the analogous case on the right side, the upper bound is taken to be infinite.
@@
For each individual power spectrum, the accretion time is determined within this interval, with the procedure applied separately for the $r$ and $\phi$ directions. The accretion time of GL-1 is estimated to be $5.6\pm 0.1$ Gyr from the $r$ direction and $5.7_{-0.4}^{+0.2}$ Gyr from the $\phi$ direction. The resulting combined accretion time is $5.6\pm 0.1$ Gyr. The quoted uncertainties reflect only the contribution from the observational errors in the data.
@@
We determine the uncertainty threshold, the bin width $\Delta$, and the bin edge shifts by a 4D grid search to maximize the significance. The uncertainty threshold and $\Delta$ are taken from $e^{-2}\approx 0.14$ to $e\approx 2.72$, with steps of 0.1 dex in log space. The bin edge shifts along each histogram dimension are taken as 0, 0.2, 0.4, 0.6 and 0.8 times the bin width. We only consider results with more than 20 stars after the uncertainty selection. Additional systematic uncertainties may arise from analysis choices such as the bin width and slit size. Moreover, systematic uncertainties introduced by the choice of potential are typically more significant.
IV.3 Effect of different potential
@@
Figure 7: Circular velocity curves $v_{c}$ of the Eilers et al. (2019) and Bovy (2015) potentials as a function of Galactocentric radius $R$.
@@
The estimated accretion times vary with the choice of the Milky Way potential model. For comparison, we adopt the potential of Bovy (2015, , hereafter B15). The circular velocity curves of this potential and the Eilers et al. (2019, , hereafter E19) potential used above are shown in Figure 7. As shown in this figure, the B15 potential has a smaller mass than the E19. At the solar position, the circular velocity is 219 km s-1 for the B15 potential and 229 km s-1 for the E19 potential.
@@
The bin width $\Delta$ and the uncertainty threshold are adjusted to maximize the significance. The right panels of Figure 5 show the 1-D power spectra of GL-1 computed using the B15 potential. It can be seen that the significances are noticeably reduced, especially for the $r$ direction. This reduction occurs when the adopted Milky Way potential poorly represents the true potential (McMillan and Binney, 2008). Nevertheless, the estimated accretion times remain close, suggesting that an incorrect potential is unlikely to fully obscure the true accretion time.
@@
In contrast to the two static potentials used in our analysis, the true potential of the Milky Way is time-dependent. The time-independent approximation is no longer reliable after $2\sim 3$ Gyr of orbital integration (Arora et al., 2022). Gómez and Helmi (2010) found in numerical simulations that accretion times inferred using time-independent approximations are underestimated by roughly 15-25% relative to the true accretion times, the exact bias depending on the speed of the potential’s evolution. This is because, as galaxies grow in mass over time, orbits at a given radius have shorter periods (higher orbital frequencies). Consequently, the spacing between the clumps in orbital frequency space used to estimate the accretion time also increases.
IV.4 Accretion times of other substructures
@@
For GL-4, the accretion time inferred from the $\phi$ direction analysis under the B15 potential is $6.9\pm 0.3$ Gyr, with a significance of $3.7\sigma$. For the E19 potential, the accretion time is estimated to be $3.7\pm 0.2$ Gyr, with a lower significance of $3.2\sigma$. The accretion time of the latter is approximately half that of the former, whereas the bin width $\Delta$ used in the analysis is approximately twice as large. This may be attributed to larger systematic uncertainties in the orbital frequencies of the E19 potential, which necessitate the use of a bin width twice as large to absorb the impact of these uncertainties. However, increasing the bin width by a factor of two degrades the resolution in orbital frequency space. As a result, the fundamental wavenumber becomes unresolvable, and only the component at half the fundamental wavenumber can be identified. As the accretion time scales linearly with the wavenumber, the inferred accretion time is therefore reduced by a factor of two.
Figure 9: The differences in orbital frequencies between the Eilers et al. (2019) and the Bovy (2015) potentials ($\Omega_{\mathrm{Bovy}}-\Omega_{\mathrm{Eilers}}$) for member stars of different substructures versus orbital pericenters under the Eilers et al. (2019) potential. Different substructures are indicated by different colors and markers. Left and right panels show the frequency differences in the $r$ and $\phi$ directions, respectively.
The larger systematic uncertainties in the orbital frequencies under the E19 potential likely stem from the smaller pericenters of the GL-4 member stars, where the two potentials differ more. Given that results based on the B15 potential show higher significance, the E19 potential therefore appears to be more strongly offset from the true potential in this region. Figure 9 shows the distribution of the differences in orbital frequencies between the two potentials, $\Delta_{\Omega}=\Omega_{\mathrm{Bovy}}-\Omega_{\mathrm{Eilers}}$, versus the orbital pericenters in the E19 potential. Member stars of different substructures are shown in different colors and markers. A negative $\Delta\Omega$ indicates that the orbital frequency is smaller in the B15 potential.
@@
Furthermore, this difference can also provide guidance for improving the potential. Eilers et al. (2019) adopted the same baryonic components as Pouliasis et al. (2017) and fitted the dark matter halo using the circular velocity curve within $5\leqslant R\leqslant 25$ kpc. The region $R<5$ kpc lies outside the fitted domain, and therefore the E19 potential is unreliable there. Since the result based on the B15 potential shows higher significance, the E19 potential may overestimate the mass within $R\sim 5$ kpc. Although the data used to fit the B15 potential also provide poor constraints in the inner region $R<5$ kpc, the B15-based result shows higher significance, suggesting that the E19 potential may overestimate the mass within $R\sim 5$ kpc.
A similar situation occurs for GR-1, as shown in Figure 9, the pericenters of its member stars are closer to the Milky Way center compared to those of GL-4, which is also the region where the E19 and B15 potentials exhibit the largest discrepancy. Its 1-D power spectra are shown in Figure 10. The results from the two potentials are $1.9\pm 0.1$ and $2.0\pm 0.1$ Gyr in the $r$ direction, which are fairly close. For the E19 potential, the significance is $3.2\sigma$, whereas for the B15 potential, it is higher, at $4.4\sigma$.
For the E19 potential, the power spectrum in the $\phi$ direction shows a trough near 1.9 Gyr. The inconsistency of the E19 potential results between the two directions suggests a limitation of the potential. Under the B15 potential, a small peak appears in the $\phi$ direction power spectrum near 2.0 Gyr, which hints that the E19 potential may overestimate the rotation curve for $R<5$ kpc. This is consistent with the inference based on the GL-4 results. However, the significance of the results based on the B15 potential is also low, indicating that the $\phi$ direction is more sensitive to the accuracy of the potential model.
The pericenters of the member stars of GR-1 are very close to the Milky Way center, where the potential is dominated by the bulge. The comparison between the E19 and B15 potential results may suggest a less massive bulge, i.e., with a mass smaller than the $1.1\times 10^{10}$ $M_{\odot}$ of the bulge in the E19 potential. For the B15 potential with a bulge mass of $0.5\times 10^{10}$ $M_{\odot}$, the peak around 2.0 Gyr in the $\phi$ direction is also very low, indicating that the bulge is likewise questionable.
@@
For the substructure GL-1, the accretion time results based on the orbital frequencies in the $r$ and $\phi$ directions show good consistency, yielding $5.6\pm 0.1$ Gyr and $5.7_{-0.4}^{+0.2}$ Gyr respectively. The statistical significance is also high, at $4.3\sigma$ and $3.9\sigma$, respectively. The agreement between the two directions and their high significance indicate the robustness of the results. The accretion time for GL-1, obtained by combining the $r$ and $\phi$ direction estimates, is $5.6\pm 0.1$ Gyr. When the B15 potential is used instead of the E19 potential, which is based on the more reliable Gaia data, the significances decrease substantially, indicating that the results are sensitive to the choice of potential.
For GL-4 the inferred accretion time is $6.9\pm 0.3$ Gyr with a significance of $3.7\sigma$, while for GR-1 it is $2.0\pm 0.1$ Gyr with a significance of $4.4\sigma$. High significance is observed in only a single direction, and the significance is greater under the B15 potential. This may be because their member stars have pericenters closer to the Milky Way center, where the E19 potential is extrapolated beyond its fitted range and thus overestimates the mass at $R<5$ kpc. For the GR-1, which lies closer to the Milky Way center, the comparison of the $r$ and $\phi$ direction estimates under the E19 potential suggests that the bulge mass may be smaller than the E19 value of $1.1\times 10^{10}$ $M_{\odot}$. For further refining accretion times for these and other phase-space substructures to obtain the accretion history of the Milky Way, more precise astrometric data provided by, e.g., Gaia DR4, China Space Station Survey Telescope and Roman Space Telescope, are needed.
All uncertainties reported for the accretion times reflect only statistical errors. The additional systematic effects arising from the adopted potential, both from the choice of the potential and from its intrinsic time evolution, are more substantial and should be addressed in future studies.
We thank the anonymous reviewer for the constructive comments and suggestions, which helped improve the clarity and quality of this paper. This work is supported by the National Key Research and Development Program of China No. 2024YFA1611902, National Natural Science Foundation of China (NSFC) No. 12588202, CAS Project for Young Scientists in Basic Research grant No. YSBR-062, the Strategic Priority Research Program of Chinese Academy of Sciences grant No. XDB1160102, and the science research grants from the China Manned Space Project with NO. CMS-CSST-2025-A11. M.C. acknowledges support in part from JSPS KAKENHI (No. JP24K00669 and 25H00394). Supported by International Partnership Program of Chinese Academy of Sciences. Grant No. 178GJHZ2022040GC. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.